# Enumerating Finite Braid Group Orbits on $SL_{2}(\mathbb{C})$-Character Varieties

### v2

@@ 1 Introduction
Dynamics of the natural braid group action on character varieties has been studied for over a century. The growing interest in classifying the finite orbits of this action has led to a wealth of new techniques and examples. Inspired by Boalch [2] who explored finite complex reflection groups as a source of finite orbits, Lam-Landesman-Litt [20] showed, up to mild conditions, all finite orbits come from finite complex reflection groups via Katz’s middle convolution operation [18]. With the goal of completing their analysis, we search the finite complex reflection groups to enumerate all possibilities and study them. We classify all finite orbits arising via middle convolution from the finite complex reflection groups.
@@
Let $\Sigma_{0,T+1}$ denote the $T$+1-punctured sphere. Given $T$ and $n$, a rank $n$ representation of the fundamental group of $\Sigma_{0,T+1}$ can be identified with a tuple of matrices $(A_{1},\dots A_{T+1})\in GL_{n}(\mathbb{C})^{T+1}$ satisfying
$\prod_{i=1}^{T+1}A_{i}=\text{Id}$
@@
The mapping class group of $\Sigma_{0,T+1}$, denoted $\operatorname{Mod}_{0,T+1}$ has a well-defined action on such representations. It is well known that $\operatorname{Mod}_{0,T+1}$ is isomorphic to the spherical braid group on $T+1$ strands, denoted $B_{T+1}$. This group has $T$ generators $\sigma_{1},\dots,\sigma_{T}$ and acts on $T$+1-tuples in the following way:
$\sigma_{i}(A_{1},A_{2},\dots,A_{T},A_{T+1})=(A_{1},\dots,A_{i-1},A_{i}A_{i+1}A_{i}^{-1},A_{i},A_{i+2},\dots,A_{T+1})$
If we further consider such representations up to simultaneous conjugation, this action descends to an action on the character variety i.e. on the isomorphism classes of representations.
@@
Conjugacy classes of tuples $[A_{1},\cdots,A_{T+1}]$ with finite orbit under this action are called MCG-finite and we say a tuple is irreducible if the associated representation is irreducible.
Recently, Lam-Landesman-Litt [20] classified finite orbits of rank two tuples with at least one matrix of infinite order. They proved that such Zariski dense representations come from pullbacks of curves or arise via middle convolution from finite complex reflection groups.
The first kind, “pullback” representations were classified by Diarra [11].
Finite irreducible complex reflection groups, which can be viewed as generalizations of Coxeter groups, were classified by Shephard and Todd in 1954 as belonging to an infinite family $G(m,p,n)$ for each rank $n$ or being one of 34 exceptional cases denoted $G_{3}$ to $G_{37}$ [9]. In view of this classification, Lam-Landesman-Litt [20, Cor. 5.1.4] prove that for $T+1\geq 7$, MCG-finite representations do not come from groups of rank five or greater. Their paper leaves open for analysis the infinite family of rank three and four and most of the exceptional groups of rank three and four, as well as the case when $T+1<7$, which we investigate.
Katz [18] originally introduced middle convolution in the language of perverse sheaves and one of our contributions is implementing computer code to explicitly compute the middle convolution given a tuple of matrices, following the algebraic reformulation of Dettweiler-Reiter [9, Section 2]. We also provide code for computing the braid group orbit for MCG-finite 4-tuples and 5-tuples in the $SL_{2}(\mathbb{C})$-character variety of the punctured sphere. This computer code is publicly accessible online at the following link: amalvayalin.github.io/research.html and we hope it encourages further applications of this operation.
@@
A $\mathbf{nice}$ $T$-tuple is a tuple (up to simultaneous conjugation) from a finite complex reflection group such that its middle convolution produces a rank 2 MCG-finite $T+1$-tuple.
For the precise definition of when a tuple is nice, see Definition 3.0.1. An explanation for how we get MCG-finite $T+1$-tuples from a $T$-tuple is provided in Remark 3.0.2. Since middle convolution depends on a parameter $\lambda$ (see Section 2.1), the same nice $T$-tuple can give rise to distinct conjugacy classes of rank 2 MCG-finite $T+1$-tuples, as discussed in Remark 3.0.3.
With this definition in mind, our problem of studying rank 2 MCG-finite tuples changes to studying nice tuples in the finite complex reflection groups. Using properties of complex reflection groups, as well as prior work on this topic, it remains to study nice $n$ and $n+1$-tuples in the rank $n=3,4$ primitive groups, nice $n+1$-tuples in the imprimitive groups of rank $2<n\leq 5$ and nice $n$-tuples in imprimitive groups of rank $n\geq 3$. More details on why these are the only remaining cases is provided in Remark 3.0.4, Remark 4.0.5 and [20, Cor. 5.1.4].
In the imprimitive groups, we either prove no such nice tuples exist, or provide a system of equations for an infinite family of solutions. We found nice tuples in all the primitive reflection groups of rank three. Among the rank four primitive reflection groups, all but $G_{29}$ and $G_{31}$ contain nice tuples. For each nice tuple in the primitive groups, we provide representatives (up to simultaneous conjugation and braid group orbit equivalence) before and after middle convolution. We also list the size of the subgroup generated by the rank 2 matrices, and in most cases we provide an approximate ‘size’ for the braid group orbit.
@@
In the rank 3 imprimitive groups, an infinite family of nice $3$-tuples exist in $G(m,1,3)$ and $G(m,m,3)$ and in each case we provide conditions on the eigenvalues of the product of the matrices in the tuple. Nice $4$-tuples exist in $G(m,p,3)$ for $m>1$ if and only if one of the following holds:
@@
In the rank 4 imprimitive groups, nice $4$-tuples and nice $5$-tuples exist in $G(m,p,4)$ if and only if $p=m$ and $m=2$ or $m=4$.
The proof of these theorems, as well as further conditions on the eigenvalues of the product of the matrices and equations for how to construct such examples are provided in Section 4.
In our computations for the primitive groups, nice tuples within each group are partitioned into different $\mathbf{types}$ based on eigenvalues (see Definition 3.2.1 and Remark 3.2.2).
@@
For each primitive reflection group of rank three, we summarize the distinct rank 2 MCG-finite tuples (up to simultaneous conjugation and braid group equivalence) produced by taking middle convolution of the different nice tuples (up to type equivalence). We also note how many of the rank 2 tuples generate an infinite subgroup.
  1. 1.
    In $G_{23}$, there are three nice 3-tuples, resulting in six distinct MCG-finite 4-tuples after middle convolution. Three of the six generate infinite subgroups. This group also has one nice 4-tuple, which after middle convolution produces one MCG-finite 5-tuple that generates an infinite subgroup.
  2. 2.
    In $G_{24}$, there is only one nice 3-tuple, which produced three distinct MCG-finite 4-tuples after middle convolution, all generating infinite subgroups. There does not exist any nice 4-tuples in this group.
  3. 3.
    In $G_{25}$, there are four nice 3-tuples, resulting in eight distinct MCG-finite 4-tuples after middle convolution. Five of these generate infinite subgroups. This group also has three nice 4-tuples, which produced two distinct MCG-finite 4-tuples after middle convolution. Only one of these 4-tuples generates an infinite subgroup.
  4. 4.
    In $G_{26}$, there are three nice 3-tuples, resulting in seven distinct MCG-finite 4-tuples after middle convolution. Five of these generate an infinite subgroup. In this group there are no nice 4-tuples.
  5. 5.
    In $G_{27}$, there are three nice 3-tuples, resulting in nine distinct MCG-finite 4-tuples after middle convolution. Six of these generate an infinite subgroup. This group also has one nice 4-tuple, which produced one MCG-finite 5-tuple after middle convolution. This 5-tuple does not generate an infinite subgroup.
@@
For each primitive reflection group of rank four, we summarize the distinct rank 2 MCG-finite tuples (up to simultaneous conjugation and braid group equivalence) produced by taking middle convolution of the different nice tuples (up to type equivalence). Only nice tuples in $G_{32}$ produces an MCG-finite tuple that generates an infinite subgroup after middle convolution.
  1. 1.
    In $G_{28}$, there is only one nice 4-tuple, which produces one MCG-finite 5-tuple after middle convolution.
  2. 2.
    In $G_{29}$ there are no nice 4-tuples nor any nice 5-tuples.
  3. 3.
    In $G_{30}$, there are three nice 4-tuples, each producing one distinct MCG-finite 5-tuple after middle convolution. For each of these 5-tuples, after tensoring with a character so the matrices are in $SL_{2}(\mathbb{C})$ we end up with one matrix being identity, which means these tuples reduce to MCG-finite 4-tuples instead. There are no nice 5-tuples in this group.
  4. 4.
    In $G_{31}$ there are no nice 4-tuples nor any nice 5-tuples.
  5. 5.
    In $G_{32}$ there are three nice 4-tuples, resulting in four distinct MCG-finite 5-tuples after middle convolution. One of these 5-tuples generates an infinite subgroup. After tensoring with a character so the matrices are in $SL_{2}(\mathbb{C})$, for two of the remaining three 5-tuples, we get the identity matrix, so these reduce to MCG-finite 4-tuples instead. This group also contains at least two nice 5-tuples, producing at least one example of an MCG-finite 6-tuple. The 6-tuple did not generate an infinite subgroup. The search for nice 5-tuples in this group was too large computationally, so this list may be incomplete.

### v3

@@ 1 Introduction
Dynamics of the natural braid group action on character varieties have been studied for over a century. The growing interest in classifying the finite orbits of this action has led to a wealth of new techniques and examples. Inspired by Boalch, [boalch2005klein] who explored finite complex reflection groups as a source of finite orbits, Lam-Landesman-Litt [LLL] showed, up to mild conditions, that all finite orbits come from finite complex reflection groups via Katz’s middle convolution operation [katz1996]. With the goal of completing their analysis, we search the finite complex reflection groups to enumerate all possibilities and study them. We classify all finite orbits arising via middle convolution from the finite complex reflection groups.
We now set notation and briefly recall the connection between local systems and representations. An expert should feel free to skip to Section 1.1.
Let $\Sigma_{T+1}$ denote the TT+1-punctured sphere, or equivalently, let $\Sigma_{T+1}=\mathbb{C}\mathbb{P}^{1}\setminus\{x_{1},\dots,x_{T+1}\}$. The fundamental group of the projective line minus $T+1$ points based at $x$ has the following presentation:
$\pi_{1}(\Sigma_{T+1},x)=\langle\gamma_{1},\dots,\gamma_{T+1}\mid\prod_{i=1}^{T+1}\gamma_{i}=1\rangle$
where each $\gamma_{i}$ is the simple closed loop around the $i$-th puncture.
A representation of $\pi_{1}(\Sigma_{T+1},x)$ is a homomorphism
$\rho:\pi_{1}(\Sigma_{T+1},x)\longrightarrow GL_{n}(\mathbb{C})$
@@
which, in the presentation given above, is completely determined by the image of each $\gamma_{i}$. Thus, a rank $n$ representation of $\pi_{1}(\Sigma_{T+1},x)$ can be identified with a tuple of matrices $(A_{1},\dots,A_{T+1})\in GL_{n}(\mathbb{C})^{T+1}$ satisfying
$\prod_{i=1}^{T+1}A_{i}=I$
where each $A_{i}=\rho(\gamma_{i})$. This condition tells us that $A_{T+1}$ is completely determined by the first $T$ matrices of the tuple. Hence, any representation of the free group on $T$ generators, denoted $F_{T}$, defined by a tuple of $T$ matrices $(A_{1},\dots,A_{T})$ gives us a representation of $\pi_{1}(\Sigma_{T+1},x)$ by defining $A_{T+1}=(\prod_{i=1}^{T}A_{i})^{-1}$ and vice versa.
For fundamental groups, changing the base-point from $x$ to $y$ is equivalent to conjugating by a path between $x$ and $y$. Thus, if we instead consider tuples of matrices up to simultaneous conjugation by an element of $GL_{n}(\mathbb{C})$, we can view these as representations of $\pi_{1}(\Sigma_{T+1})$, without specifying a base point. On representations, this corresponds to the action of $GL_{n}(\mathbb{C})$ via
$g\cdot\rho=g\rho g^{-1}.$
Quotienting by this action gives us isomorphism classes of representations, and the resulting space is called the character variety (See Section 2.3). For tuples of matrices $(A_{1},\dots,A_{T+1})$, we write $A_{1},\dots,A_{T+1}]$ for the corresponding equivalence class in the character variety.
A local system $\mathcal{L}$ with fiber isomorphic to a vector space $V$ on $\Sigma_{T+1}$ is a sheaf on $\Sigma_{T+1}$ locally isomorphic to the constant sheaf defined by $V$. At each point $x\in\Sigma_{T+1}$, we have an isomorphism of the stalk $\mathcal{L}_{x}\cong V$; the dimension of $V$ (as a vector space over $\mathbb{C}$) is called the rank of the local system $\mathcal{L}$. For each loop $\gamma:[0,1]\longrightarrow\Sigma_{T+1}$ with $\gamma(0)=\gamma(1)$, we have isomorphisms
$V\cong\mathcal{L}_{\gamma(0)}\cong\mathcal{L}_{\gamma(1)}\cong V.$
The isomorphism $V\rightarrow V$ given by composing the above maps only depends on the homotopy class of $\gamma$ and corresponds to some invertible linear transformation $A\in GL(V)$. We call $A$ the monodromy of $\mathcal{L}$ along $\gamma$.
Given a local system $\mathcal{L}$ of rank $n$ on $\Sigma_{T+1}$, the monodromy along each generator $\gamma_{i}$ of $\pi_{1}(\Sigma_{T+1},x)$ provides an element $A_{i}$ of $GL_{n}(\mathbb{C})$, and these $A_{i}$ are referred to as the local monodromy. The representation defined by $\rho(\gamma_{i})=A_{i}$ is called the monodromy representation, and its image (as a subgroup of $GL_{n}$) is called the monodromy group. Moreover, given a representation $\rho:\pi_{1}(\Sigma_{T+1},x)\longrightarrow GL_{n}(\mathbb{C})$, a local system $\mathcal{L}$ with monodromy determined by $\rho$ can be constructed via the universal cover of $\Sigma_{T+1}$. We refer the reader to [locsysexpo] for more exposition on local systems.
The above correspondence between local systems and representations is well-behaved on isomorphism classes, so we get the following 3-way bijections:
[…]
$\qquad\qquad\qquad\updownarrow$
$\left\{\text{Conjugacy classes of representations $\pi_{1}(\Sigma_{T+1})\xrightarrow{\rho}GL_{n}(\mathbb{C})$}\right\}$
$\qquad\qquad\qquad\updownarrow$
[…]
1.1 Mapping Class Group and Middle Convolution
On the character variety, we also have the following operations:
  1. 1.
    Action of the mapping class group: Associated to the surface $\Sigma_{T+1}$ is its mapping class group
    $\operatorname{Mod}_{T+1}=\pi_{0}(\operatorname{Homeo}^{+}(\Sigma_{T+1}))$
    which has an action on representations of $\pi_{1}(\Sigma_{T+1})$ via the natural action on $\Sigma_{T+1}$. Note that since $\operatorname{Mod}_{T+1}$ typically does not preserve the base point of the fundamental group, we really do want to consider conjugacy classes of representations.
@@ 1.1 Mapping Class Group and Middle Convolution
    It is well known that $\operatorname{Mod}_{T+1}$ is isomorphic to the spherical braid group on $T+1$ strands, denoted $B_{T+1}$. This group has $T$ generators $\sigma_{1},\dots,\sigma_{T}$ and acts on TT+1-tuples in the following way:
    $\sigma_{i}([A_{1},A_{2},\dots,A_{T},A_{T+1}])=[A_{1},\dots,A_{i-1},A_{i}A_{i+1}A_{i}^{-1},A_{i},A_{i+2},\dots,A_{T+1}].$
    By definition of this action, we see that $\sigma_{i}([A_{1},A_{2},\dots,A_{T},A_{T+1}])$ will always be a tuple consisting of matrices from the group $\langle A_{1},\dots,A_{T+1}\rangle$. Hence, whenever the $A_{i}$ all live in some finite group, the corresponding tuple will have finite orbit under this braid group action.
  2. 2.
    Middle convolution: Let $\operatorname{LocSys}(\Sigma_{T+1})$ denote isomorphism classes of local systems on $\Sigma_{T+1}$ for any rank $n$. For each $\lambda\in\mathbb{C}^{\times}\setminus\{1\}$ , Katz [katz1996] introduces an invertible functor called middle convolution:
    $MC_{\lambda}:\operatorname{LocSys}(\Sigma_{T+1})\longrightarrow\operatorname{LocSys}(\Sigma_{T+1}).$
    This operation takes an isomorphism class of rank $n$ local systems on $\Sigma_{T+1}$ to an isomorphism class of rank $n^{\prime}$ local systems on $\Sigma_{T+1}$. The rank $n^{\prime}$ depends on $\lambda$ and on the conjugacy class of the local monodromy. For a more precise definition of middle convolution of local systems, see Lam-Landesman-Litt[LLL, Definition 3.1.1]. For tuples of matrices, the corresponding operation of middle convolution takes as input only the first $T$ matrices of the tuple and is discussed further in Section 2.1. This algebraic version of middle convolution, formulated for representations of free group on $T$ generators, was introduced by Dettweiler-Reiter in [KatzAlgo, Prop. 2.6] and [MC-L], and is mainly what we use in this paper. Properties relevant to our work are that middle convolution is equivariant with respect to the action of the braid group [KatzAlgo, Thm. 5.1, Cor. 3.6] and preserves irreducibility, which we define below.
@@
Representations, or equivalently tuples $A_{1},\cdots,A_{T+1}]$ with $\prod_{i=1}^{T+1}A_{i}=I$, in the character variety with finite orbit under the mapping class group action are called MCG-finite. We say a tuple is irreducible if the associated representation is irreducible.
Recently, Lam-Landesman-Litt [LLL] classified MCG-finite tuples of rank two with at least one matrix of infinite order. They proved that such Zariski-dense representations come from pullbacks of curves or arise via middle convolution of local systems whose monodromy group is a finite complex reflection group.
The first kind, “pullback” representations, were classified by Diarra [diarra2013, Sections 3-5].
Finite irreducible complex reflection groups, which can be viewed as generalizations of Coxeter groups, were classified by Shephard and Todd in 1954 as belonging to an infinite family $G(m,p,n)$ (imprimitive groups) for each rank $n$ or being one of 34 exceptional cases (primitive groups) denoted $G_{3}$ to $G_{37}$ [stfinite]. Each of these groups $W$ comes with a canonical finite-dimensional complex representation (given by reflection matrices acting on some vector space $V$), and the rank of $W$ is defined as the rank of this representation. The rank of $W$ will be the rank of the local system taken as input for middle convolution.
In view of this classification, Lam-Landesman-Litt [LLL, Cor. 5.1.4] prove that for $T+1\geq 7$, MCG-finite representations do not come from reflection groups of rank five or greater. Their paper leaves open for analysis the imprimitive groups of rank three and four and most of the primitive groups of rank three and four, as well as the case when $T+1<7$, which we investigate.
One wants to classify remaining cases of MCG-finite rank two local systems on $\Sigma_{T+1}$, knowing that they come by middle convolution applied to certain rank $n$ local systems on $\Sigma_{T+1}$ with monodromy group a finite complex reflection group $W$. As a result, the issue is to classify those rank $n$ local systems on $\Sigma_{T+1}$ (or equivalently, $T+1$-tuples of $n\times n$ matrices multiplying to the identity that come from a rank $n$ complex reflection group $W$), which indeed give rise by middle convolution to a rank two local system on $\Sigma_{T+1}$ with finite MCG-orbit.
Using that $T+1$-tuples which multiply to the identity are determined by the first $T$ matrices in the tuple, we restrict our attention to $T$-tuples in reflection groups with desirable properties.
@@
A $T$-tuple $\textbf{A}=[A_{1},A_{2},\dots,A_{T}]$ of matrices (defined up to simultaneous conjugation) from a finite complex reflection group $W$ is called $\mathbf{nice}$ if A satisfies the following:
  1. 1.
    All $A_{i}$ are reflections in $W$
  2. 2.
    $W=\langle A_{1},A_{2},\dots,A_{T}\rangle$
  3. 3.
    $(A_{1}A_{2}\cdots A_{T})^{-1}$ has an eigenvalue $\lambda\neq 1$ of multiplicity T2T-2.
Note that by Condition 3, nice tuples should never multiply to the identity, which is why we restrict to only the first $T$ matrices of the $T+1$-tuple corresponding to our rank $n$ local system on $\Sigma_{T+1}$. The reason for each of these conditions is explained in more detail in Section 3, but we briefly outline the idea here.
Remark 1.1.3.
Given a nice $T$-tuple $\mathbf{A}=[A_{1},A_{2},\dots,A_{T}]$, any $\lambda$ satisfying Condition 3 ensures that $MC_{\lambda}(\mathbf{A})=[\tilde{A_{1}},\tilde{A_{2}},\dots,\tilde{A_{T}}$] is a $T$-tuple of rank two (see Section 2.1). Taking the $T+1$-th matrix to be
$\tilde{A}_{T+1}=(\Pi_{i=1}^{T}\tilde{A_{i}})^{-1}$
[…]
We have written computer code for each step of the process described above. We start by searching finite complex reflection groups for nice tuples. Since middle convolution depends on a parameter $\lambda$, the same nice $T$-tuple can be used to produce non-isomorphic rank 2 local systems on $\Sigma_{T+1}$, as discussed in Remark 1.1.8. Although Katz [katz1996] originally introduced middle convolution in the language of perverse sheaves, one of our contributions is implementing computer code to explicitly compute the middle convolution given a tuple of matrices and parameter $\lambda\in\mathbb{C}^{\times}$, following the algebraic reformulation of Dettweiler-Reiter [MC-L, Section 2]. We also provide code for computing the braid group orbit for MCG-finite tuples in the $SL_{2}(\mathbb{C})$-character variety of $\Sigma_{T+1}$, for $T=3,4$. All the computer code is publicly accessible online at the following link: amalvayalin.github.io/research.html, and we hope it encourages further applications of this operation.
We now provide a list of exclusions that this work will not treat:
  1. i.
    We will not consider local systems on $\Sigma_{T+1}$ produced from local systems on $\Sigma_{T}$ by inserting the identity matrix into the $T$-tuple. This method ‘artificially’ produces MCG-finite local systems on the sphere with more punctures, so we consider only tuples without any identity matrix.
  2. ii.
    We will not consider nice $T$-tuples $A_{1},\dots,A_{T}]$ from finite complex reflection groups such that their inverse product $A_{T+1}=(A_{1}\cdots A_{T})^{-1}$ is a scalar matrix $\lambda^{-1}I\in GL_{n}$. Writing
    $MC_{\lambda}(A_{1},\dots,A_{T})=[\tilde{A_{1}},\dots,\tilde{A_{T}}$], we have a local system of rank 2. Because of how eigenvalues transform under middle convolution (see Theorem 2.1.6), we know that $\tilde{A}_{T+1}=(\prod_{i=1}^{T}\tilde{A_{i}})^{-1}$ will be the scalar $\lambda I\in GL_{2}$. Multiplying by a character to put this tuple into $SL_{2}$, we still have that the new $T+1$-th matrix $\tilde{A^{\prime}}_{T+1}$ of the tuple is a scalar matrix. In $SL_{2}$, the only scalar matrices are $I$ or I-I. If $\tilde{A^{\prime}}_{T+1}=I$, then we exclude it by the previous point above. If $\tilde{A^{\prime}}_{T+1}=-I$, then for some non-scalar $\tilde{A^{\prime}_{j}}$, $j\neq T+1$, we can construct an equivalent tuple by replacing $\tilde{A^{\prime}_{j}}$ with $\tilde{A^{\prime}_{j}}$ and replacing $\tilde{A^{\prime}}_{T+1}$ by $\tilde{A^{\prime}}_{T+1}$. But this means $\tilde{A^{\prime}}_{T+1}=I$, and this tuple will be excluded.
  3. iii.
    We do not search finite reflection groups of rank 2 for nice tuples. These can only produce local systems on $\Sigma_{3}$ or $\Sigma_{4}$ and we would be going from a rank 2 local system to another rank 2 local system, possibly with infinite monodromy group after middle convolution. For $\Sigma_{3}$, the action of the braid group factors through $S_{3}$ (see [farbmarg, Prop. 2.3]), so all orbits are MCG-finite. For $\Sigma_{4}$, rank 2 local systems over the 4-punctured sphere have been studied in connection to the Painlevé VI equation (see Section 1.2) and finite orbits were classified by Lisovyy-Tykhyy[T-L].
  4. iv.
    We do not consider non-irreducible tuples. These are classified by Cousin-Moussard [cousinmouss, Section 3].
With the above exclusions in mind, we discuss which cases remain to be considered.
One property of finite complex reflection groups (see Section 2.2) is that a group of rank $n$ is generated either by $n$ reflections or $n+1$ reflections. By the definition of nice tuples, the condition of generating the whole group Condition 2 tells us we need $n\leq T$, and since an $n\times n$ matrix cannot have an eigenvalue of multiplicity greater than $n$, the eigenvalue condition Condition 3 tells us $T-2\leq n$. However, if T2=nT-2=n, then $A_{T+1}$ will be a scalar matrix (as Jordan blocks with nonzero eigenvalue of size >1>1 do not have finite order), hence falls under Exclusion ii, so in fact we require T2<nT-2<n, or equivalently $T<n+2$. Thus, for rank 2 local systems on $\Sigma_{T+1}$, we need only search for nice $T$-tuples in reflection groups of rank $n$ satisfying the relation
$n\leq T\leq n+1.$
[…]
Thus, we search for nice tuples in finite complex reflection groups $W$ with $\operatorname{rank}(W)\in\{3,4,5\}$ with hopes of using middle convolution to produce rank 2 MCG-finite local systems on $\Sigma_{4},\Sigma_{5}$, and $\Sigma_{6}$.
We note now that the only rank 5 primitive group is $G_{33}$, also denoted as $W(K_{5})$, which Lam-Landesman-Litt [LLL, Lemma 5.4.1] show cannot produce an irreducible rank 2 MCG-finite local system via middle convolution.
For imprimitive groups $W=G(m,p,n)$, the parameter $n$ denotes the rank of $W$. The case where $n\geq 5$ is easiest to consider, as these groups do not contain any nice $n$-tuples.
@@
The next two theorems pertain to the imprimitive groups of rank $3$ and $4$. In these groups, we either prove no nice tuples exist or provide a system of equations for an infinite family of solutions. Section 4 provides the proofs of the following theorems, as well as conditions on the eigenvalues of the product of the matrices in the tuple, or equivalently, conditions on the parameter $\lambda$ that will be used for middle convolution whenever nice tuples are shown to exist.
@@
In the rank 3 imprimitive groups, there are no nice 3-tuples in $G(m,p,3)$ if $p\notin\{1,m\}$, and an infinite family of nice $3$-tuples exists in $G(m,1,3)$ and $G(m,m,3)$ for all $m>1$. Nice $4$-tuples exist in $W=G(m,p,3)$ for $m>1$ if and only if $W$ is one of the following groups:
@@
In the rank 4 imprimitive groups, nice $4$-tuples and nice $5$-tuples exist in $W=G(m,p,4)$ if and only if $W$ is one of the following groups:
  1. 1.
    $G(2,2,4)$
  2. 2.
    $G(4,4,4)$
The next collection of theorems all pertain to primitive reflection groups of rank 3 and 4. For these groups, we performed a computer search for nice tuples and analyzed our findings. In our computations, nice tuples within each group were partitioned into different $\mathbf{types}$ based on eigenvalues (see Definition 3.2.1 and Remark 3.2.2) and we counted the nice tuples in each group by type. We summarize the distinct rank 2 MCG-finite tuples (up to simultaneous conjugation and braid group equivalence) produced by taking middle convolution of the different nice tuples (up to type equivalence). We also note whether any of the tuples generate an infinite subgroup of $SL_{2}$, providing examples of MCG-finite local systems with infinite monodromy group.
@@
All primitive reflection groups of rank 3 contain nice 3-tuples, and all but $G_{24}$ and $G_{26}$ contain nice 4-tuples. Table 2 and Table 2 summarize the data before and after applying middle convolution.
Summary of MCG-finite Rank 2 Local Systems on $\Sigma_{4}$ Produced from Rank 3 Primitive Groups
Rank 3 Primitive Group # Nice 3-Tuples # MCG-finite Rank 2 Tuples aaaaaaaaaaaaaaa # Tuples Generating Infinite Subgroup of $SL_{2}(\mathbb{C})$
$G_{23}$ 3 6 3
$G_{24}$ 1 3 3
$G_{25}$ 4 8 5
$G_{26}$ 3 7 5
$G_{27}$ 3 9 6
Table 1:
Summary of MCG-finite Rank 2 Local Systems on $\Sigma_{5}$ Produced from Rank 3 Primitive Groups
Rank 3 Primitive Group # Nice 4-Tuples # MCG-finite Rank 2 Tuples aaaaaaaaaaaaaaa # Tuples Generating Infinite Subgroup of $SL_{2}(\mathbb{C})$
$G_{23}$ 1 1 1
$G_{24}$ NA - -
$G_{25}$ 3 2 1
$G_{26}$ NA - -
$G_{27}$ 1 1 0
Table 2:
Remark 1.1.8.
For some nice $T$-tuples, it is possible that several eigenvalues $\lambda$ have multiplicity T2T-2 (Condition 3). Hence, the different choices of parameter $\lambda$ give rise to possibly different rank 2 MCG-finite $T+1$-tuples after middle convolution. For example, when $T=3$, we only need eigenvalues of multiplicity one, which is why each type of nice 3-tuple can produce up to three distinct MCG-finite 4-tuples. It is also possible that different eigenvalues or different types produce equivalent MCG-finite tuples after middle convolution, resulting in fewer MCG-finite tuples than nice tuples. An example of this is the nice 4-tuples produced from the rank 3 group $G_{25}$.
Now we discuss primitive groups of rank four.
@@
In $G_{28}$, there is only one nice 4-tuple, which produces one MCG-finite 5-tuple after middle convolution. There are no nice 5-tuples in this group.
Theorem 1.1.10.
In $G_{29}$ and $G_{31}$, there are no nice tuples.
For the remaining two primitive groups of rank 4, some of the nice tuples exhibit interesting behaviour. After applying middle convolution, when we put the matrices into $SL_{2}(\mathbb{C})$, we have that one of the matrices of the MCG-finite TT+1-tuple becomes the identity (Exclusion i). This means they come from MCG-finite local systems on $\Sigma_{T}$ (i.e. MCG-finite $T$-tuples) by the operation of inserting the identity matrix as the TT+1-th matrix. In this way, we get that our nice $T$-tuples produce MCG-finite $T$-tuples of rank 2 rather than the expected MCG-finite TT+1-tuples of rank 2.
Theorem 1.1.11.
In $G_{30}$, there are three nice 4-tuples, each producing one distinct MCG-finite 5-tuple after middle convolution. All three of these 5-tuples comes from 4-tuples by inserting the identity, so the nice 4-tuples in this group produce MCG-finite 4-tuples and no MCG-finite 5-tuples. There are no nice 5-tuples in this group.
Theorem 1.1.12.
In $G_{32}$, there are three nice 4-tuples, resulting in four distinct MCG-finite 5-tuples after middle convolution. One of these 5-tuples generates an infinite subgroup. Two of the remaining three 5-tuples come from MCG-finite 4-tuples by inserting the identity matrix. This group also contains at least two nice 5-tuples, producing at least one example of an MCG-finite 6-tuple. The 6-tuple does not generate an infinite subgroup. In total, $G_{32}$ produced two MCG-finite 4-tuples, two MCG-finite 5-tuples, and at least one MCG-finite 6-tuple.
Note that the search for nice 5-tuples in $G_{32}$ was too large computationally, so this list may be incomplete. For each nice tuple in the primitive groups, we provide representatives (up to simultaneous conjugation and braid group orbit equivalence) before (Section 5) and after middle convolution (in Appendix A, Appendix B, Appendix C). We also list the size of the subgroup generated by the rank 2 matrices (after middle convolution), and in most cases we provide an approximate ‘size’ for the braid group orbit in the appendices.
A consolidated summary of which finite complex reflection groups contain nice tuples that can be used to produce MCG-finite rank 2 local systems on $\Sigma_{T+1}$ is given in Table 4 and Table 4.
Rank($W$) aaaaaaaaaaaaaaa Imprimitive Groups that Product MCG-finite Rank 2 Local Systems on $\Sigma_{4}$ aaaaaaaaaaaaaaa Imprimitive Groups that Product MCG-finite Rank 2 Local Systems on $\Sigma_{5}$ aaaaaaaaaaaaaaa Imprimitive Groups that Product MCG-finite Rank 2 Local Systems on $\Sigma_{6}$
3 $G(m,1,3),G(m,m,3)$ a $G(m,m,3)$, for $m\neq 3$, $G(m,1,3)$, for $\gcd(m,3)=1$ $G(m,3,3)$, $m\neq 3$ NA
4 NA $G(2,2,4),G(4,4,4)$ $G(2,2,4),G(4,4,4)$
$\geq 5$ NA NA NA
Table 3: For all the above, we assume $m>1$.
Finite Complex Reflection Groups aaaaaaaaaaaaaaa aaaaaaaaaaaaaaa #MCG-finite Rank 2 Local Systems Produced on $\Sigma_{4}$ aaaaaaaaaaaaaaa #MCG-finite Rank 2 Local Systems Produced on $\Sigma_{5}$ aaaaaaaaaaaaaaa #MCG-finite Rank 2 Local Systems Produced on $\Sigma_{6}$
$G_{23}$ 6 1 0
$G_{24}$ 3 0 0
$G_{25}$ 8 2 0
$G_{26}$ 7 0 0
$G_{27}$ 9 1 0
$G_{28}$ 0 1 0
$G_{29}$ 0 0 0
$G_{30}$ 3 0 0
$G_{31}$ 0 0 0
$G_{32}$ 2 2 At least 1
Table 4: